# Video2BEV: Transforming Drone Videos to BEVs for Video-based Geo-localization

### v1

@@ 4.1 Video2BEV Transformation
Figure 4: The overview of the Video2BEV paradigm. Video2BEV Transformation (left). Given drone-view video $V_{drn}$ containing multi-view frames, we adopt 3D Gaussian Splatting (3DGS) to reconstruct the scene at first. Then we render the scene from a Bird-Eye-View to get the projection (BEVs). Considering the region of the core area, we further crop BEVs for training. We can observe that BEVs exhibit resemblances to the corresponding satellite-view images. Hard Negative Sample Synthesis (right). Given captions generated by an off-the-shelf visual-LLM [11], we fine-tune a stable-diffusion model [28] with LoRA [10], and conduct inference to synthesize samples which serve as negative samples for subsequent usage. Model Architecture (middle). Given outputs of the proposed Video2BEV transformation, we extract embeddings by a shared encoder for satellite images $x_{sat}$ and BEVs $x_{bev}$, supervised by the contrastive loss $\mathcal{L}_{C}$ and the instance loss $\mathcal{L}_{I}$. Then we extract embeddings from synthetic BEVs $x_{s\text{-}bev}$ and adopt Multilayer Perceptrons (MLPs) to fuse both positive and negative samples, supervised by the matching loss $\mathcal{L}_{M}$. Similar operations are also applied for satellite-view images which are omitted.
During the flight around the target building, the viewpoints of the camera vary, resulting in captured drone-view videos that contain rich multi-view information about both the target building and surrounding areas. Rather than taking viewpoint variation as an obstacle for aligning satellite and drone views, we explicitly leverage the multi-view information and transform the drone-view video into Bird’s Eye Views (BEVs). In doing so, we ease the learning process for the model. Instead of learning geometry correspondence and feature correspondences simultaneously, the model only needs to learn the feature mapping relationship between two views, thus significantly facilitating network convergence. Illustrated in the left of Fig. 4, given the drone-view video containing multi-view images, we estimate corresponding camera poses by structure from motion [39] and reconstruct the scene containing the target buildings utilizing 3D Gaussian Splatting (3DGS) [12]. After reconstructing the scene, we adopt the normalized input pose and the unit vector in the world coordinate to calculate the BEV camera pose and render BEVs. For the training set, we incorporate rotation angles and varying heights into the BEV camera pose, generating a sequence of rotating and scaled-down BEVs. In terms of the test set, we only increase the height of the BEV camera poses and render a sequence of scaled-down BEV images. The number of BEVs corresponds to the number of images in the drone video. As shown in Fig. 4, outputs of Video2BEV transformation do not suffer from severe distortion, thereby aiding in the establishment of fine-grained spatial correspondence with the satellite view.
@@ 4.2 Hard Negative Sample Synthesis
Negative samples play a significant role in discriminative metric learning. Current negative sample mining strategies cannot ensure the quantity and quality of negative samples, as the number of challenging samples is limited, and selected negative samples can not always share similar architectural styles and other details with the original samples. In order to bypass these drawbacks, we propose to fine-tune a diffusion model to generate diverse samples as negative samples, shown in the right part of Fig. 4. After transforming the drone-view video to BEVs $x_{bev}$ via the Video2BEV transformation, we utilize a visual-LLM [11] to generate captions in a multi-round process for both BEV and satellite-view images, during which a human annotator interacts with visual-LLM to restrict the token length of the captions. After obtaining captions for the BEV and satellite images, we fine-tune a stable diffusion network [28] with LoRA [10] to generate diverse synthetic images, during which we freeze the CLIP text encoder [25]. The outputs of this model are challenging negative samples for the subsequent step. Considering the transformed BEV and satellite images share the same viewing direction and content, we conduct inference on the same stable diffusion network to generate negative samples for BEV and satellite-view images by adopting corresponding captions. We provide visualizations of original and synthetic images in Fig. 5. Similar to the original samples, synthetic negative samples exhibit comparable appearances, including color patterns and structural features of buildings.
@@ 4.3 Model Optimization
In this section, we do not overstate the novelty of the architecture. Instead, we adopt a general architecture from vision-language models [14, 15], enhanced by the synthetic negative samples. The model architecture is shown in the middle of Fig. 4 and is optimized in a two-stage manner. In the first stage, we transform the drone-view video to BEVs by the proposed Video2BEV transformation (see the left part of Fig. 4). Then, we adopt a shared encoder to extract embeddings from paired BEV and satellite-view images. The encoder is ViT-S [6] excluding the classifier. The supervisions of this stage are the instance loss $\mathcal{L}_{I}$ with the square-ring partition [42] and the contrastive loss $\mathcal{L}_{C}$ [14]. For the instance loss, we first partition the embeddings from different views into different parts adopting the square-ring partition [42] strategy. Next, we apply multiple classifier modules to each part of the embeddings (similar to LPN [42]), yielding the location probability. Then we accumulate instance losses from multiple parts, and the instance loss $\mathcal{L}_{I}$ [55] is formulated as the location classification:
$\mathcal{L}_{I}=-log({\hat{p}_{sat}})-log({\hat{p}_{bev}}),$ (1)
where $\hat{p}_{sat}$ and $\hat{p}_{bev}$ are the predicted probability that belongs to the ground-truth label from two views respectively. For the contrastive loss, given a pair of satellite-view and BEV images, the satellite-to-BEV similarity is defined as:
@@
where $f_{sat}$ and $f_{bev}$ are the embeddings of the same location from two platforms, and $f^{j}_{bev}$ denotes the sample within the mini-batch. $\tau$ is a learnable temperature parameter. $s(\cdot,\cdot)$ denotes the cosine similarity. Similarly, the BEV-to-satellite similarity is $S_{bev2sat}$ and the contrastive loss $\mathcal{L}_{C}$ is:
@@
For the second stage, we employ two-layer Multilayer Perceptrons (MLPs) alongside the square-ring partition [42] to fuse two embeddings derived from different views of the input including BEV, satellite, and synthetic BEV, and synthetic satellite view images. Then we concatenate all parts of embeddings from two views and project the fused embeddings into the two-dimensional space by another MLPs. The matching loss of two-view inputs $\mathcal{L}_{M}$ is:
@@
where $\hat{p}_{m}$ is the estimated matching probability and $p_{m}$ is a ground-truth binary label. If the two input data do not contain synthetic data and are from the same category, then $p_{m}=1$; otherwise, $p_{m}=0$. Specifically, for the BEVs, we calculate the matching loss two times. For the first calculation, we rank the similarity $S$ and select three negative samples from the satellite-view images, ensuring that these samples do not belong to the same category. For the second calculation, we similarly select another three negative samples from the synthetic BEV images, which includes samples from the same category that are actually challenging negative samples. We apply similar operations for the satellite-view input as well. Finally, we accumulate and average matching losses across different combinations of inputs. In summary, the loss functions in our method include the instance loss $\mathcal{L}_{I}$, the contrastive loss $\mathcal{L}_{C}$, and the matching loss $\mathcal{L}_{M}$. Specifically, we train the first stage of our method (including the encoder, classifier modules, and the temperature parameter $\tau$) with the instance loss $\mathcal{L}_{I}$ and the contrastive loss $\mathcal{L}_{C}$ at first. Subsequently, we freeze the fine-tuned first-stage weights and train the second stage (including two types of MLPs) from scratch, supervised by the matching loss $\mathcal{L}_{M}$.
Discussion. What are the advantages of the synthetic negative samples? First, inspired similar success in other fine-grained tasks [52, 33, 53], we encourage the model “see” more samples to prevent over-fitting as well as facilitate discriminative feature learning. With the assistance of the diffusion model, we can synthesize negative samples of a diverse range of categories and quantities. In terms of categories, the synthetic samples exhibit similar yet different architectural styles as shown in  Fig. 5, including variations in color and pattern, as well as structural details such as the shape and material. Regarding quantities, the negative sample pool is no longer constrained to a fixed size. Utilizing the diffusion model, we can theoretically generate an infinite number of images as negative samples, expanding the negative sample pool significantly. Second, there is a clear relationship between our synthetic negative samples and the anchor samples. This is because we use identical captions from the original samples to synthesize the negative samples. This relationship ensures that our negative samples are appropriately challenging.
@@ 5 Experiment
Implementation Details. For the synthetic negative samples, the captions for BEV and satellite images differ and we synthesize 32 negative samples for both BEV and satellite images of each category. We train the first stage of the proposed model with the AdamW optimizer, with a batch size of 140, for 140 epochs, and a learning rate of $2e^{-5}$ and $2e^{-4}$ for the encoder and other modules in the first stage respectively. Then we freeze parameters in the first stage and train the second stage from scratch with a similar training configuration. During the test stage, we utilize the similarity scores from the first stage of the proposed model to select the top 32 samples from the gallery, and then apply the second stage of the proposed model to re-rank these top 32 samples. The whole framework is implemented with Pytorch.
Evaluation Metrics. Satellite-view data is in image format, while drone-view data is collected in video format. We can treat drone view data as images or video. In this paper, we adopt the video setting for the evaluation of competitive methods and our method. Specifically, we treat a drone video as an individual query or gallery by averaging the similarity scores of the images within the video in a late fusion manner. In our method, there is a similar averaging operation on similarity scores of the BEVs which is also in video format. We employ 2-fps videos in UniV as the training and test sets and release videos at 5 and 10 fps for further research usage.
Table 2: Comparisons on the UniV dataset for geo-localization between Drone (D) and Satellite (S). All methods are compared in the video setting. R@1 is recall at top1. AP (%) is average precision (high is good). $\theta$ denotes the elevation angle of the drone flight. The proposed method yields the best results.
@@
Figure 7: Qualitative results for Drone $\rightarrow$ Satellite and Satellite $\rightarrow$ Drone. We replace the visualization of the drone-view videos with the transformed output (BEVs) as the query or gallery. Given queries (left) from different platforms, matched galleries are in green box, and mismatched galleries are in blue box. The scores on the top are similarity scores estimated by the proposed method.
5.1 Comparisons with Competitive Methods
Quantitative Results. As shown in Tab. 2, we compare the proposed method with other competitive methods on the UniV dataset. The performance of our method has surpassed that of other competitive methods [42, 3, 43, 5]. On the $45^{\circ}$ subset, our method achieves gains of 0.30% Recall and 0.58% AP for satellite $\rightarrow$ drone compared to the second-best method. On the $30^{\circ}$ subset, all methods experience a performance drop. As shown in Fig. 6, we highlight some imperfect reconstructed regions by the Video2BEV transformation. The lower elevation of the drone flights raises more occlusions (see Tab. 1), which also compromises our Video2BEV transformation. Compared to the second best method, our method is still robust, receiving improvements of 3.2% AP for drone $\rightarrow$ satellite and 5.8% AP for satellite $\rightarrow$ drone respectively (see Tab. 2). All methods are compared in the video setting, which means we temporally average the outputs of frames in a video from the drone view. For methods with officially released weights (Sample4Geo, DWDR), we test these methods on the $45^{\circ}$ test set directly and subsequently retrain and evaluate these methods on the $30^{\circ}$ subsets. For methods without official weights (LPN, FSRA), we retrain these methods on both the $45^{\circ}$ and $30^{\circ}$ subsets to ensure a fair comparison.
Qualitative Results. We show qualitative results of the drone geo-localization in Fig. 7. In our method, drone-view videos are transformed to BEVs by the proposed Video2BEV transformation and we choose the representative sample from the BEV sequence for visualization. For drone $\rightarrow$ satellite, we observe that the proposed method effectively retrieves reasonable buildings with similar structural features, such as cross-shaped roofs and roofs equipped with solar panels. For satellite $\rightarrow$ drone, we find a similar result. Our method successfully retrieves true-matched results at the top of the candidate list among images with similar contents. We add more qualitative visualizations including failure case analysis in the supplementary material.
@@ 5.2 Ablation Studies and Further Discussion
Effect of Training Strategies. We explore three different strategies for training the proposed model. For the Train Together strategy, we load the matched weights pre-trained on the ImageNet dataset [27]. Then we fine-tune the first stage of the proposed model and train the second stage of the proposed model from scratch simultaneously. The Fine-tune strategy entails loading fine-tuned weights of the first stage on the UniV dataset. After this, we fine-tune the first stage with a smaller learning rate while simultaneously training the second stage from scratch. The Freeze strategy consists of loading fine-tuned weights of the first stage on UniV, then fixing all weights of the first stage, while training the second stage from scratch at the same time. The results of three training strategies are in Tab. 3(b). The Train Together strategy yields the worst results. We attribute this to the difficulty of training both stages simultaneously, as the first stage of the proposed model is designed for coarse-grained retrieval, while the second stage of the proposed model focuses on fine-grained retrieval, which relies on the output of the first stage. When both stages are trained together, the first stage fails to retrieve reliable candidates for the second stage, affecting the overall training process. The Fine-tune strategy achieves a significant performance boost, as the first stage is able to produce reliable embeddings for the second stage. Finally, we freeze the first stage after loading its corresponding weight. The Freeze strategy receives the best result, and we adopt this training strategy.
Effect of Re-ranking Top-K Samples. During the test stage, we select top-k samples from the gallery leveraging the similarity score from the first stage of our method and subsequently re-rank these samples by the second stage. We conduct hyper-parameter experiments with varying values of top-k. As shown in Tab. 3(c), we select k $\in$ $\{8,16,32,64,128,256,512\}$. Re-ranking the top-512 and top-256 samples yields the best performance and re-ranking the top-256, top-128, top-64, and top-32 samples results in a slight performance drop respectively. Re-ranking the top-16 and top-8 samples leads to a further decline in performance. Considering the balance between the performance of our method and the testing time, we choose to re-rank the top 32 samples as default.

### v3

@@ 4.1 Video2BEV Transformation
Figure 4: The overview of the Video2BEV paradigm. Video2BEV Transformation (left). Given drone-view video $V_{drn}$ containing multi-view frames, we adopt 3D Gaussian Splatting (3DGS) to reconstruct the scene at first. Then we render the scene from a Bird-Eye-View to get the projection (BEVs). Considering the region of the core area, we further crop BEVs for training. We can observe that BEVs exhibit resemblances to the corresponding satellite-view images. Hard Negative Sample Synthesis (right). Given captions generated by an off-the-shelf visual-LLM [16], we fine-tune a stable-diffusion model [39] with LoRA [15], and conduct inference to synthesize samples which serve as negative samples for subsequent usage. Model Architecture (middle). Given outputs of the proposed Video2BEV transformation, we extract embeddings by a shared encoder for satellite images $x_{sat}$ and BEVs $x_{bev}$, supervised by the contrastive loss $\mathcal{L}_{C}$ and the instance loss $\mathcal{L}_{I}$. Then we extract embeddings from synthetic BEVs $x_{s\text{-}bev}$ and adopt MLP to fuse both positive and negative samples, supervised by the matching loss $\mathcal{L}_{M}$. Similar operations for satellite-view images are omitted.
@@
During the flight around the target location, the viewpoints of the camera change dynamically, resulting in captured drone-view videos that contain rich multi-view information about both the target location and surrounding areas. We explicitly leverage the multi-view information and transform the drone-view video into Bird’s Eye Views (BEVs). In doing so, we ease the learning process for the subsequent model. Instead of learning geometry correspondence and feature correspondences simultaneously [4], the subsequent model only needs to learn the feature mapping relationship between two views, thus significantly facilitating network convergence. As shown in the left part of Fig. 4, given the drone-view video containing multi-view images, we estimate corresponding camera poses by structure from motion [53] and reconstruct the scene containing the target locations utilizing 3D Gaussian Splatting (3DGS) [18]. After reconstructing the scene, we adopt the normalized camera pose and the unit vector in the world coordinate to calculate the BEV camera pose and render BEVs. In particular, the vanilla 3DGS takes less than 8 seconds to render 50 BEVs with the shape of 512 $\times$ 512 on a NVIDIA 4090 GPU. Since we usually foreknow the search area in practice, the BEV generation process can be regarded as an off-line pre-processing. BEVs rendered with our Video2BEV transformation module do not suffer from severe distortion (see  Fig. 4), thereby aiding in the subsequent inter-platform matching process with the satellite images.
@@ 4.2 Hard Negative Sample Synthesis
Negative samples play a significant role in discriminative metric learning. Current negative sample mining strategies [8, 31, 71] cannot ensure the quantity and quality of negative samples, as challenging samples are inherently scarce, and the selected negative samples do not necessarily exhibit similar architectural styles or consistent semantic details as the original samples. In order to bypass these drawbacks, we propose to generate diverse BEV representations as hard negative samples through a fine-tuned diffusion model, which is shown in the right part of Fig. 4. After transforming the drone-view video to BEVs $x_{bev}$ via our Video2BEV transformation, we utilize a visual-LLM [16] to generate captions in a multi-round manner for both BEV and satellite-view images. After obtaining captions for the BEVs and satellite images, we fine-tune a stable diffusion network [39] with LoRA [15] to generate diverse synthetic images, during which we freeze the CLIP text encoder [35]. The outputs of this model are negative samples for the subsequent step. Since the transformed BEV and satellite images share the same top-down viewpoint and semantic content, we conduct inference on the same diffusion model to generate negative samples for both BEV and satellite-view images using their corresponding captions. We provide visualizations of original and synthetic images in Fig. 5. Synthetic negative samples exhibit similar semantic contents as the original ones but with different fine-grained information, thus enhancing the discriminative intra-platform representation learning.
@@ 4.3 Model Optimization
We adopt a general architecture from vision-language models [24, 25], enhanced by BEVs and synthetic negative samples. The model architecture is shown in the middle of Fig. 4 and is optimized in a two-stage manner following [25, 68]. In the first stage, we transform the drone-view video to BEVs by the proposed Video2BEV transformation (see the left part of Fig. 4). Then, we adopt a shared encoder to extract embeddings from paired BEV and satellite-view images. The encoder is ViT-S [10] excluding the classifier. The supervisions of this stage are the instance loss $\mathcal{L}_{I}$ [75] with the square-ring partition [58] and the contrastive loss $\mathcal{L}_{C}$ [24]. We apply multiple classifier modules to each part of the embeddings (similar to LPN [58]), yielding the location probability of two views which are denoted as $\hat{p}_{sat}$ and $\hat{p}_{bev}$ respectively. The instance loss $\mathcal{L}_{I}$ is formulated as the location classification as :
$\mathcal{L}_{I}=-log({\hat{p}_{sat}})-log({\hat{p}_{bev}}).$ (1)
Then we accumulate instance losses from multiple parts to form the final instance loss. For the contrastive loss, given a pair of satellite-view and BEV images, the satellite-to-BEV similarity is defined as:
@@
where $f_{sat}$ and $f_{bev}$ are embeddings of the same location from two platforms, and $f^{j}_{bev}$ denotes the sample within the mini-batch. $\tau$ is a learnable temperature parameter. $s(\cdot,\cdot)$ denotes the cosine similarity. Similarly, the BEV-to-satellite similarity is $S_{bev2sat}$ and the contrastive loss $\mathcal{L}_{C}$ is:
@@
In the second stage, we employ a two-layer MLP alongside the square-ring partition [58] to fuse two embeddings obtained from anchor-positive or anchor-negative samples Specifically, when BEV serves as the anchor, satellite and synthetic BEV serve as the positive sample and the negative sample, respectively. Similarly, when satellite acts as positive samples, BEV and synthetic satellite act as positive and negative samples respectively. Then we project the fused embeddings into the two-dimensional space using another MLP. Given inputs from paired samples, the matching loss $\mathcal{L}_{M}$ between them is calculated as:
@@
where $\hat{p}_{m}$ is the estimated matching probability and $p_{m}$ is a ground-truth binary label. If the two input data do not contain the synthetic data and are both from the same location, then $p_{m}=1$; otherwise, $p_{m}=0$. Specifically, for the BEVs, we calculate the matching loss two times. For the first calculation, we rank the similarity $S$ and select three negative samples from the satellite-view images, ensuring that these samples do not belong to the same location simultaneously. For the second calculation, we similarly select another three negative samples from the synthetic BEVs, which are actually hard negatives from the same location. We apply similar operations to the satellite-view input. Finally, we accumulate and average matching losses across different combinations of inputs. In summary, the loss functions in our method include the instance loss $\mathcal{L}_{I}$, the contrastive loss $\mathcal{L}_{C}$, and the matching loss $\mathcal{L}_{M}$. Specifically, in the first stage of training, we optimize the encoder, classifier modules of our model, the temperature parameter $\tau$ with the instance loss $\mathcal{L}_{I}$ and the contrastive loss $\mathcal{L}_{C}$. Subsequently, we freeze the parameters fine-tuned in the first stage and train the MLPs from scratch in the second stage under the supervision of matching loss $\mathcal{L}_{M}$.
Discussion. What are the advantages of the synthetic negative samples? Inspired by successes in other fine-grained tasks [72, 44, 73], we encourage the model “see” more samples to prevent over-fitting as well as facilitate discriminative intra-platform feature learning. We are similar to GeNIe [20] in that both methods alter the image representation of the target object to generate hard negative samples. However, it is worth noting that there are two primary differences. (1) We have a larger modification space, and the negative sample pool is no longer constrained to a fixed size. Different from the GeNIe in changing limited categories for classification, we perturb the initial noise of the diffusion model, and it leads to diverse generations. Utilizing the diffusion model, we can theoretically generate an infinite number of images as negative samples, expanding the negative sample pool significantly. (2) We retain the semantic content of the anchor samples in our synthetic negative samples. This is because we employ identical captions from the original samples to synthesize the negative samples while only changing the initial noise. This slight modification ensures that our negative samples are appropriately challenging, and encourages the model to check the fine-grained discrepancies among semantical-similar samples.
@@ 5 Experiment
Implementation Details. Since the captions for the two views (drone and satellite) employ different wording to describe the same location, we synthesize samples based on the text and generate separate negative samples for each view. All input images are resized to 256$\times$256. We train the first stage of the proposed model with the AdamW optimizer, with a batch size of 140, for 140 epochs, and a learning rate of $2e^{-5}$ and $2e^{-4}$ for the encoder and other modules in the first stage respectively. Then we freeze parameters in the first stage and train the second stage from scratch with a similar training configuration. During the test stage, we utilize the similarity scores from the first stage of ours to select the top 32 samples from the gallery, and then apply the second stage of the proposed model to re-rank these top 32 samples. The framework is implemented with PyTorch.
Evaluation Metrics. Satellite-view data is in image format, while drone-view data is collected in video format. We can treat drone view data as images or video. In this paper, we adopt the video setting for the evaluation of competitive methods and our method. Specifically, we treat a drone video as an individual query or gallery by averaging the similarity scores of the images within the video in a late fusion manner. There is a similar averaging operation on similarity scores of the BEVs which is also in video format.
Table 2: (a) Comparisons on the UniV for geo-localization between Drone (D) and Satellite (S) platforms. R@1 is recall at top1. AP (%) is average precision (high is good). (b) Comparisons in terms of an out-of-distribution testing on the SUES-200 ($45^{\circ}$ test set). Our method still yields the best results.
@@
5.1 Comparisons with Competitive Methods
Quantitative Results. As shown in Tab. 2(a), we compare the proposed method with other competitive methods on the UniV dataset. The performance of our method has surpassed that of other competitive methods [58, 6, 59, 8]. On the $45^{\circ}$ subset, our method achieves gains of 0.30% Recall and 0.58% AP for satellite $\rightarrow$ drone compared to the second-best method. On the $30^{\circ}$ subset, all methods experience a performance drop. As shown in Fig. 6, we highlight some imperfect reconstructed regions by the Video2BEV transformation. The lower elevation of the drone flights raises more occlusions (see Tab. 1), which also compromises our Video2BEV transformation. Compared to the second best method, our method is still robust, receiving improvements of 3.2% AP for drone $\rightarrow$ satellite and 5.8% AP for satellite $\rightarrow$ drone respectively (see Tab. 2(a)). All methods are compared in the video setting, which means we temporally average the outputs of frames in a video from the drone view. For methods with officially released weights (Sample4Geo [8], DWDR [59]), we test these methods on the $45^{\circ}$ test set directly and subsequently retrain and evaluate these methods on the $30^{\circ}$ subsets. For methods without official weights (LPN [58], FSRA [6]), we retrain them on both the $45^{\circ}$ and $30^{\circ}$ subsets to ensure a fair comparison.
Out-of-Distribution (OOD) Scalability. We also evaluate the model trained on the UniV dataset ($45^{\circ}$) on the unseen SUES-200 $45^{\circ}$ test set in an OOD manner. SUES-200 dataset contains dense frames collected in real-world environments, including real-world light, shadow transformations, and disturbances. We observe that our method shows strong OOD potential, surpassing the runner-up method by more than 7% AP for drone $\rightarrow$ satellite (see Tab. 2(b)).
Qualitative Results. We show qualitative results of the drone geo-localization on the UniV and SUES-200 datasets (see Fig. 7). In our method, drone-view videos are transformed to BEVs by the proposed Video2BEV transformation and we choose the representative sample from the BEV sequence for visualizations. For drone $\rightarrow$ satellite, we observe that the proposed method effectively retrieves reasonable locations with similar structural features, such as cross-shaped roofs and roofs equipped with solar panels. For satellite $\rightarrow$ drone, we find a similar result. Our method successfully retrieves true-matched results at the top of the candidate list among images with similar contents. We add more visualizations in the supplementary material.
@@ 5.1 Comparisons with Competitive Methods
Figure 7: Qualitative results on the UniV (a, c) and SUES-200 (b,d) dataset for Drone $\rightarrow$ Satellite and Satellite $\rightarrow$ Drone. We depict the transformed output (BEVs) as the query or gallery. Given queries (left) from different platforms, matched galleries are in green box, and mismatched galleries are in blue box. The scores on the top are similarity scores estimated by the proposed method.
@@ 5.2 Ablation Studies and Further Discussion
Effect of Training Strategies. We explore three different strategies for training. For the Train Together strategy, we load the matched weights pre-trained on the ImageNet dataset [38]. Then we fine-tune the first stage of the proposed model and train the second stage of the proposed model from scratch. The Fine-tune strategy entails loading fine-tuned weights of the first stage on the UniV dataset. After this, we fine-tune the first stage with a smaller learning rate while training the second stage from scratch. The Freeze strategy consists of loading fine-tuned weights of the first stage on UniV, then fixing all weights of the first stage, while training the second stage from scratch. The results of three training strategies are in Tab. 3(b). The Train Together strategy yields the worst results. We attribute this to the difficulty of training both stages simultaneously, as the first stage of the proposed model is designed for coarse-grained retrieval, while the second stage of the proposed model focuses on fine-grained retrieval, relying on the output of the first stage. When both stages are trained together, the first stage fails to retrieve reliable candidates for the second stage, affecting the overall training process. The Fine-tune strategy achieves a significant performance boost, as the first stage is able to produce reliable embeddings for the second stage. Finally, we freeze the first stage after loading its corresponding weight. The Freeze strategy yields the best result, and we adopt this strategy.
Effect of Re-ranking Top-K Samples. During the test stage, we select top-k samples from the gallery leveraging the similarity score from the first stage of our method and subsequently re-rank these samples by the second stage. We conduct hyper-parameter experiments with varying values of top-k, and select k $\in$ $\{8,16,32,64,128,256,512\}$ (see  Tab. 3(c)). Re-ranking the top-512 and top-256 samples yields the best performance and re-ranking the top-256, top-128, top-64, and top-32 samples results in a slight performance drop respectively. Re-ranking the top-16 and top-8 samples leads to a further decline in performance. Considering the balance between the performance and the testing time, we re-rank the top 32 samples as default.